OpenStudy (anonymous):

Use the Law of Sines to find the missing angle of the triangle. Find measure of angle b given that c=83 , a=44 and that measure of angle a =31 . A=76.3 degrees B= 15.8 degrees C= 72.7 degrees

OpenStudy (anonymous):

Substitute these values in the Law of Sines and you will get your answer

OpenStudy (anonymous):

don't really understand laws of sines or anything that's why I posted it to get help

OpenStudy (anonymous):

visit and watch this video http://www.youtube.com/watch?v=5WJKmirJWjA

jimthompson5910 (jim_thompson5910):

you say that the measure of angle a is 31, but you also say that A = 76.3 degrees

jimthompson5910 (jim_thompson5910):

which one is it?

OpenStudy (anonymous):

the capital a is the choice a b or c sorry

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the whole problem?

jimthompson5910 (jim_thompson5910):

there are too many a's, b's and c's...lol

OpenStudy (anonymous):

here are the choice s A=76.3 degrees B= 15.8 degrees C= 72.7 degrees heres the question its confusing and I cant do screen shots sorry Use the Law of Sines to find the missing angle of the triangle. Find measure of angle b given that c=83 , a=44 and that measure of angle a =31

jimthompson5910 (jim_thompson5910):

ok so let me get this straight...choice A is 76.3, choice B is 15.8 and choice C is 72.7 right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

the law of sines is sin(A)/a = sin(C)/c plug in the given info to get sin(31)/44 = sin(C)/83 now solve for angle C sin(31)/44 = sin(C)/83 83*sin(31)/44 = sin(C) 0.97154909585 = sin(C) ... use a calculator here sin(C) = 0.97154909585 C = arcsin(0.97154909585) C = 76.2999941 ... use a calculator here C = 76.3 Now use this to find angle B A+B+C = 180 31 + B + 76.3 = 180 B + 107.3 = 180 B = 180 - 107.3 B = 72.7 So angle B is 72.7 degrees

OpenStudy (anonymous):

okay thank you for the help everyone

jimthompson5910 (jim_thompson5910):

which is choice C again...too many A's, B's and C's

OpenStudy (anonymous):

thank you for your help though thanks a lot

jimthompson5910 (jim_thompson5910):

you're welcome